# Hallo4: High-Fidelity Dynamic Portrait Animation via Direct Preference Optimization and Temporal Motion Modulation

### v1

@@ 1 Introduction
Portrait image animation, an interdisciplinary challenge at the intersection of computer vision and computer graphics, aims to synthesize photorealistic 2D videos of humans driven by multimodal control signals such as text, audio, and pose. This technology holds significant potential for applications spanning digital entertainment, virtual reality, human-computer interaction, and personalized marketing. Despite recent advances in diffusion models [9, 11, 43], vision transformers [22, 15, 23], and autoregressive architectures [32, 2, 16], achieving high-fidelity animation remains constrained by two persistent challenges: (1) generating lip-sync and facial expressions that are both perceptually natural and aligned with human preferences, and (2) capturing high-frequency motions, such as subtle articulations, dynamic expressions, and rapid body gestures—particularly during abrupt speech or sudden hand movements. In this work, we specifically investigate audio-driven animation of facial and upper-body portraits enhanced by skeletal motion conditions, aiming to address these two key challenges.
Direct preference optimization (DPO) [44], a policy alignment framework that bypasses explicit reward modeling by optimizing generative models directly through pairwise human preference data, has gained increasing attention in text-to-image and text-to-video generation domains [34]. Unlike conventional reinforcement learning approaches, DPO minimizes distributional shifts during generation by aligning outputs with human preferences—such as aesthetic quality, semantic coherence, and temporal consistency—while eliminating the need for explicit reward model training and adversarial optimization loops through pairwise preference rankings with minimal data. Building on this paradigm, we introduce the first audio-driven portrait DPO dataset specifically designed for human-centric animation, capturing human preferences along two critical axes: (1) portrait-motion-video synchronization accuracy and (2) facial expression and pose naturalness. By contrasting trajectory probabilities of preferred and dispreferred samples using a KL-constrained reward function, our method optimizes the generation policy to maximize trajectory-level reward margins while regularizing deviations from the base diffusion model’s denoising dynamics. This approach achieves measurable improvements in lip-sync accuracy (matching real-data benchmarks in lip-sync scores) and enhances facial expressiveness, establishing a new baseline for preference-aligned portrait animation.
@@ 2 Related Work
Portrait Image Animation. Research in portrait animation has evolved through distinct technical paradigms addressing sequential limitations. Early lip-synchronization methods, exemplified by Wav2Lip [26], employed adversarial discriminators to enhance audio-visual alignment but lacked holistic facial dynamics. Subsequent 3D morphable model (3DMM)-based approaches improved motion naturalness through structured representations: SadTalker [45] disentangled audio-driven 3D coefficients, while VividTalk [28] and DreamTalk [41] introduced hybrid mesh-codebook strategies. Diffusion-driven frameworks advanced realism by integrating facial landmarks (DiffTalk [27], AniPortrait [38]) and hierarchical temporal conditioning (Loopy [14], Hallo [39]), with LivePortrait [10] optimizing efficiency via implicit keypoints. Audio-centric methods like Sonic [13] prioritized global perceptual coherence for diverse motion synthesis. Transformer-based architectures, including VASA-1 [40], Hallo3 [7] and OmniHuman [19], unified facial and bodily dynamics in latent spaces for partial and full-body generalization. Our work introduces a human-annotated DPO dataset and optimization approach to enhance lip-sync accuracy and facial naturalness, complementing existing UNet and DiT-based audio-driven diffusion frameworks.
@@
Human Preference Alignment. Research on aligning generative models with human preferences spans visual understanding and synthesis tasks. In video understanding, early efforts [29] adapted reinforcement learning with human feedback to mitigate hallucinations, while subsequent works introduced self-supervised alignment via language model rewards [44], AI feedback [1], and temporal grounding optimization [17]. For image generation, Diffusion-DPO [34] reformulated DPO to align text-to-image outputs with human judgments. Video generation methods further extended DPO to optimize visual-textual semantics (VideoDPO [21]), enable online training (OnlineVPO [42]), and refine motion dynamics (Flow-DPO [20]). While these works establish DPO as a scalable paradigm for preference-aware synthesis, none address the unique challenges of audio-driven portrait animation. To our knowledge, this work presents the first curated human preference dataset and tailored DPO framework for enhancing lip-sync accuracy and facial naturalness in portrait animation domain.
@@
Motion Condition. Strategies for motion condition injection in human video generation depend on motion signal type and control granularity. Audio-driven portrait methods, such as EMO [31], VASA-1 [40], and FantasyTalking [37], leverage cross-attention to align audio features with visual frames. Body motion control approaches like CHAMP [47], MimicMotion [46], and EasyControl [36] employ in-context learning for coarse-grained skeletal dynamics. Half-body animation frameworks, including CyberHost [18], EchoMimicV2 [25], and OmniHuman [19], combine cross-attention for facial synchronization with in-context learning for skeletal motion, achieving coherent upper-body motion. Building on these insights, we propose a temporal motion condition modulation mechanism that enhances high-frequency motion details and fine-grained articulatory synchronization, particularly optimizing motion fidelity for DiT-based portrait diffusion models.
@@ 3.2 Preference-Driven Portrait Animation
Preference Dataset Curation. Direct preference optimization frames human preference alignment as a policy optimization problem using pairwise preference data $\mathcal{D}=\{(x,y^{w},y^{l})\}$, where $y^{w}$ is preferred over $y^{l}$; here, $y^{w}$ and $y^{l}$ represent the “winning” and “losing” samples, respectively [34]. To construct this dataset, we employ annotators to assess the generated samples across the two dimensions and combine them with a “best-vs-worst” selection strategy to obtain the final preference dataset.
@@ 3.3 Unified Temporal Motion Modulation
Latent Video Representation. The original video data $\mathbf{V}$ is encoded into a compressed latent representation $\mathbf{Z}$ via a pretrained causal 3D variational autoencoder. Following established methods [35], the input video $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$ is compressed to latent features $\mathbf{Z}=\mathcal{E}(\mathbf{V})\in\mathbb{R}^{T^{\prime}\times H^{\prime}\times W^{\prime}\times d}$, where $T^{\prime}=\lfloor T/4\rfloor$ represents the temporally downsampled length, and $H^{\prime}=\lfloor H/8\rfloor$, $W^{\prime}=\lfloor W/8\rfloor$ denote spatially reduced dimensions. This encoding balances computational efficiency with the preservation of critical motion dynamics and spatiotemporal coherence inherent in the original video.
Human Motion Condition. We process two complementary motion modalities to govern facial articulations and body dynamics. For audio-driven facial movements, the raw speech signals $\mathbf{S}\in\mathbb{R}^{L}$ are encoded into frame-aligned acoustic features $\mathbf{C}_{\text{audio}}\in\mathbb{R}^{T\times N_{a}\times d_{a}}$ via Wav2Vec 2.0 [3], preserving phoneme-level temporal precision essential for lip synchronization. For skeletal-guided body motion, 2D human keypoints $\mathbf{V}_{k}\in\mathbb{R}^{T\times J\times 3}$ (with $J$ joints) are processed through an efficient convolutional network, generating motion descriptors $\mathbf{C}_{\text{skel}}\in\mathbb{R}^{T\times H^{\prime}\times W^{\prime}\times d_{k}}$ that maintain original temporal resolution while achieving spatial alignment with video latents through dimension reduction to $H^{\prime}\times W^{\prime}$. This dual-conditioning mechanism encourages synchronized integration of high-frequency facial movements and coarse body kinematics with the latent video representation.
@@ 3.4 DPO-Enhanced Diffusion Network
For the UNet architecture, we adopt the design principles of EMO [30], Hallo [39], and EchoMimic-v2 [5], using Stable Diffusion 1.5 [4] as the backbone. A parallel ReferenceNet structure preserves visual identity consistency, while audio-conditioning layers inject speech features through cross-attention mechanisms. The diffusion denoising process employs denoising diffusion probabilistic models, which learns a parameterized reverse process through variational lower bound optimization to iteratively recover clean data from noise.
For the DiT architecture, we utilize the Wan2.1 framework [35] as the backbone, consistent with FantasyTalking [37]. This architecture employs a 3D causal VAE for spatiotemporal compression and decoding. The transformer module processes video data by unfolding it into token sequences via 3D patchify operations. Each transformer block integrates self-attention for spatial relations, cross-attention for text conditioning, and timestep embeddings. All transformer layers share a unified MLP structure. The denoising process leverages flow matching principles, constructing vector fields based on optimal transport path conditional probability flows to directly regress velocity fields. This enables deterministic ODE-based transformations from noise distributions to data distributions.
@@ 3.5 Training and Inference
For UNet-based diffusion, we initialize with the pretrained Hallo portrait model [39] and fine-tune it using direct preference optimization. The DPO objective $\mathcal{L}_{\text{DPO}}$ optimizes denoising trajectories to align with human preference judgments through gradient updates on the UNet parameters.
@@ 4.1 Experimental Setups
UNet-based SadTalker [45] 7.804 7.956 11.826
DreamTalk [24] 7.515 7.717 9.142
AniPortrait [38] 3.550 11.007 14.819
Sonic [13] 8.186 6.823 9.370
Hallo [39] 7.770 7.605 8.508
@@
DiT-based Hallo3 [7] 7.384 8.613 8.589
FantasyTalking [37] 4.218 11.043 9.806
@@
UNet-based SadTalker [45] 5.126 8.219 43.499 64.919 463.671
DreamTalk [24] 4.691 8.290 68.211 137.049 879.902
AniPortrait [38] 2.129 10.911 31.760 63.106 509.498
Sonic [13] 5.387 7.370 19.928 56.104 489.458
Hallo [39] 4.880 8.189 27.252 51.009 469.985
@@
DiT-based Hallo3 [7] 4.648 8.992 23.671 54.105 397.425
FantasyTalking [37] 2.957 10.559 47.124 69.332 614.462
@@ 4.2 Experimental Setups
Evaluation Metrics. For motion-video alignment, we adopt the widely used Sync-C and Sync-D metrics [6] to measure the confidence and distance between visual and audio content. Hand quality and motion richness are measured using hand keypoint confidence (HKC) [18] and hand keypoint variance (HKV) [18], respectively. These metrics validate the DPO results after integrating preference data with human motion alignment, as well as the improvements in highly dynamic motion enabled by the unified temporal motion modulation module. To evaluate facial fidelity, we employ E-FID [8], which quantifies expression discrepancies by computing the Fréchet Inception Distance (FID) of facial expression parameters extracted via face reconstruction. For visual quality assessment, we utilize FID, FVD [33], PSNR, and SSIM.
Baseline. We conduct comprehensive comparisons across two dominant architectural paradigms for audio-driven portrait animation: UNet-based and DiT-based approaches. For UNet-based methods, we benchmark against SadTalker [45], DreamTalk [24], AniPortrait [38], Sonic [13], and Hallo [39]. Our UNet variant, denoted as Ours (UNet), extends Hallo by integrating human preference data and direct preference optimization (DPO). Within the DiT-based category, we evaluate Hallo3 [7] and FantasyTalking [37], while our DiT implementation (Ours (DiT)) builds upon the Wan2.1 backbone [35], incorporating unified temporal motion modulation and DPO-enhanced training. To validate generalization to skeletal-guided synthesis, we further compare with pose-driven half-body animation methods: AnimateAnyone [12], MimicMotion [46], and EchoMimic-v2 [25].
@@ 4.3 Comparison with State-of-the-Art
AnimateAnyone [12] - 34.350 0.824 20.482 0.761 51.439 453.528
MimicMotion [46] - 34.783 0.846 19.186 0.713 49.039 468.004
EchoMimic-v2 [25] 6.624 33.284 0.863 21.421 0.796 38.461 301.347
@@ 4.4 Ablation Study and Discussion
FantasyTalking [37] 3.38 3.33 3.50 2.73
DreamTalk [24] 2.83 2.88 2.68 3.20
Hallo3 [7] 3.90 3.73 3.93 3.80
Sonic [13] 4.18 4.08 4.13 4.03

### v2

@@ 1 Introduction
Portrait image animation, an interdisciplinary challenge at the intersection of computer vision and computer graphics, aims to synthesize photorealistic 2D videos of humans driven by multimodal control signals such as text, audio, and pose. This technology holds significant potential for applications spanning digital entertainment, virtual reality, human-computer interaction, and personalized marketing. Despite recent advances in diffusion models esser2024scaling ; ho2020denoising ; zhang2023adding , vision transformers liu2024vmamba ; kirillov2023segment ; liu2022swin , and autoregressive architectures touvron2023llama ; alayrac2022flamingo ; li2023blip , achieving high-fidelity animation remains constrained by two persistent challenges: (1) generating lip-sync and facial expressions that are both perceptually natural and aligned with human preferences, and (2) capturing high-frequency motions, such as subtle articulations, dynamic expressions, and rapid body gestures—particularly during abrupt speech or sudden hand movements. In this work, we specifically investigate audio-driven animation of facial and upper-body portraits enhanced by skeletal motion conditions, aiming to address these two key challenges.
Direct preference optimization (DPO) zhang2024direct , a policy alignment framework that bypasses explicit reward modeling by optimizing generative models directly through pairwise human preference data, has gained increasing attention in text-to-image and text-to-video generation domains wallace2024diffusion . Unlike conventional reinforcement learning approaches, DPO minimizes distributional shifts during generation by aligning outputs with human preferences—such as aesthetic quality, semantic coherence, and temporal consistency—while eliminating the need for explicit reward model training and adversarial optimization loops through pairwise preference rankings with minimal data. Building on this paradigm, we introduce the first audio-driven portrait DPO dataset specifically designed for human-centric animation, capturing human preferences along two critical axes: (1) portrait-motion-video synchronization accuracy and (2) facial expression and pose naturalness. By contrasting trajectory probabilities of preferred and dispreferred samples using a KL-constrained reward function, our method optimizes the generation policy to maximize trajectory-level reward margins while regularizing deviations from the base diffusion model’s denoising dynamics. This approach achieves measurable improvements in lip-sync accuracy (matching real-data benchmarks in lip-sync scores) and enhances facial expressiveness, establishing a new baseline for preference-aligned portrait animation.
@@ 2 Related Work
Portrait Image Animation. Research in portrait animation has evolved through distinct technical paradigms addressing sequential limitations. Early lip-synchronization methods, exemplified by Wav2Lip prajwal2020lip , employed adversarial discriminators to enhance audio-visual alignment but lacked holistic facial dynamics. Subsequent 3D morphable model (3DMM)-based approaches improved motion naturalness through structured representations: SadTalker zhang2022sadtalker disentangled audio-driven 3D coefficients, while VividTalk sun2023vividtalk and DreamTalk zhang2023dream introduced hybrid mesh-codebook strategies. Diffusion-driven frameworks advanced realism by integrating facial landmarks (DiffTalk shen2023difftalk , AniPortrait wei2024aniportrait ) and hierarchical temporal conditioning (Loopy jiang2024loopy , Hallo xu2024hallo ), with LivePortrait guo2024liveportrait optimizing efficiency via implicit keypoints. Audio-centric methods like Sonic ji2024sonic prioritized global perceptual coherence for diverse motion synthesis. Transformer-based architectures, including VASA-1 xu2024vasa , Hallo3 cui2024hallo3 and OmniHuman lin2025omnihuman , unified facial and bodily dynamics in latent spaces for partial and full-body generalization. Our work introduces a human-annotated DPO dataset and optimization approach to enhance lip-sync accuracy and facial naturalness, complementing existing UNet and DiT-based audio-driven diffusion frameworks.
@@
Human Preference Alignment. Research on aligning generative models with human preferences spans visual understanding and synthesis tasks. In video understanding, early efforts sun2023aligning adapted reinforcement learning with human feedback to mitigate hallucinations, while subsequent works introduced self-supervised alignment via language model rewards zhang2024direct , AI feedback ahn2024tuning , and temporal grounding optimization li2025temporal . For image generation, Diffusion-DPO wallace2024diffusion reformulated DPO to align text-to-image outputs with human judgments. Video generation methods further extended DPO to optimize visual-textual semantics (VideoDPO liu2024videodpo ), enable online training (OnlineVPO zhang2024onlinevpo ), and refine motion dynamics (Flow-DPO liu2025improving ). While these works establish DPO as a scalable paradigm for preference-aware synthesis, none address the unique challenges of audio-driven portrait animation. To our knowledge, this work presents the first curated human preference dataset and tailored DPO framework for enhancing lip-sync accuracy and facial naturalness in portrait animation domain.
@@
Motion Condition. Strategies for motion condition injection in human video generation depend on motion signal type and control granularity. Audio-driven portrait methods, such as EMO tian2021good , VASA-1 xu2024vasa , and FantasyTalking wang2025fantasytalking , leverage cross-attention to align audio features with visual frames. Body motion control approaches like CHAMP zhu2024champ , MimicMotion zhang2024mimicmotion , and EasyControl wang2024easycontrol employ in-context learning for coarse-grained skeletal dynamics. Half-body animation frameworks, including CyberHost lin2024cyberhost , EchoMimicV2 meng2024echomimicv2 , and OmniHuman lin2025omnihuman , combine cross-attention for facial synchronization with in-context learning for skeletal motion, achieving coherent upper-body motion. Building on these insights, we propose a temporal motion condition modulation mechanism that enhances high-frequency motion details and fine-grained articulatory synchronization, particularly optimizing motion fidelity for DiT-based portrait diffusion models.
@@ 3.2 Preference-Driven Portrait Animation
Preference Dataset Curation. Direct preference optimization frames human preference alignment as a policy optimization problem using pairwise preference data $\mathcal{D}=\{(x,y^{w},y^{l})\}$, where $y^{w}$ is preferred over $y^{l}$; here, $y^{w}$ and $y^{l}$ represent the “winning” and “losing” samples, respectively wallace2024diffusion . To construct this dataset, we employ annotators to assess the generated samples across the two dimensions and combine them with a “best-vs-worst” selection strategy to obtain the final preference dataset.
@@ 3.3 Unified Temporal Motion Modulation
Latent Video Representation. The original video data $\mathbf{V}$ is encoded into a compressed latent representation $\mathbf{Z}$ via a pretrained causal 3D variational autoencoder. Following established methods wan2025 , the input video $\mathbf{V}\in\mathbb{R}^{T\times H\times W\times 3}$ is compressed to latent features $\mathbf{Z}=\mathcal{E}(\mathbf{V})\in\mathbb{R}^{T^{\prime}\times H^{\prime}\times W^{\prime}\times d}$, where $T^{\prime}=\lfloor T/4\rfloor$ represents the temporally downsampled length, and $H^{\prime}=\lfloor H/8\rfloor$, $W^{\prime}=\lfloor W/8\rfloor$ denote spatially reduced dimensions. This encoding balances computational efficiency with the preservation of critical motion dynamics and spatiotemporal coherence inherent in the original video.
Human Motion Condition. We process two complementary motion modalities to govern facial articulations and body dynamics. For audio-driven facial movements, the raw speech signals $\mathbf{S}\in\mathbb{R}^{L}$ are encoded into frame-aligned acoustic features $\mathbf{C}_{\text{audio}}\in\mathbb{R}^{T\times N_{a}\times d_{a}}$ via Wav2Vec 2.0 baevski2020wav2vec , preserving phoneme-level temporal precision essential for lip synchronization. For skeletal-guided body motion, 2D human keypoints $\mathbf{V}_{k}\in\mathbb{R}^{T\times J\times 3}$ (with $J$ joints) are processed through an efficient convolutional network, generating motion descriptors $\mathbf{C}_{\text{skel}}\in\mathbb{R}^{T\times H^{\prime}\times W^{\prime}\times d_{k}}$ that maintain original temporal resolution while achieving spatial alignment with video latents through dimension reduction to $H^{\prime}\times W^{\prime}$. This dual-conditioning mechanism encourages synchronized integration of high-frequency facial movements and coarse body kinematics with the latent video representation.
@@ 3.4 DPO-Enhanced Diffusion Network
For the UNet architecture, we adopt the design principles of EMO tian2024emo , Hallo xu2024hallo , and EchoMimic-v2 chen2024echomimic , using Stable Diffusion 1.5 blattmann2023stable as the backbone. A parallel ReferenceNet structure preserves visual identity consistency, while audio-conditioning layers inject speech features through cross-attention mechanisms. The diffusion denoising process employs denoising diffusion probabilistic models, which learns a parameterized reverse process through variational lower bound optimization to iteratively recover clean data from noise.
For the DiT architecture, we utilize the Wan2.1 framework wan2025 as the backbone, consistent with FantasyTalking wang2025fantasytalking . This architecture employs a 3D causal VAE for spatiotemporal compression and decoding. The transformer module processes video data by unfolding it into token sequences via 3D patchify operations. Each transformer block integrates self-attention for spatial relations, cross-attention for text conditioning, and timestep embeddings. All transformer layers share a unified MLP structure. The denoising process leverages flow matching principles, constructing vector fields based on optimal transport path conditional probability flows to directly regress velocity fields. This enables deterministic ODE-based transformations from noise distributions to data distributions.
@@ 3.5 Training and Inference
For UNet-based diffusion, we initialize with the pretrained Hallo portrait model xu2024hallo and fine-tune it using direct preference optimization. The DPO objective $\mathcal{L}_{\text{DPO}}$ optimizes denoising trajectories to align with human preference judgments through gradient updates on the UNet parameters.
@@ 4.1 Experimental Setups
UNet-based SadTalker zhang2022sadtalker 7.804 7.956 11.826
DreamTalk ma2023dreamtalk 7.515 7.717 9.142
AniPortrait wei2024aniportrait 3.550 11.007 14.819
Sonic ji2024sonic 8.186 6.823 9.370
Hallo xu2024hallo 7.770 7.605 8.508
@@
DiT-based Hallo3 cui2024hallo3 7.384 8.613 8.589
FantasyTalking wang2025fantasytalking 4.218 11.043 9.806
@@
UNet-based SadTalker zhang2022sadtalker 5.126 8.219 43.499 64.919 463.671
DreamTalk ma2023dreamtalk 4.691 8.290 68.211 137.049 879.902
AniPortrait wei2024aniportrait 2.129 10.911 31.760 63.106 509.498
Sonic ji2024sonic 5.387 7.370 19.928 56.104 489.458
Hallo xu2024hallo 4.880 8.189 27.252 51.009 469.985
@@
DiT-based Hallo3 cui2024hallo3 4.648 8.992 23.671 54.105 397.425
FantasyTalking wang2025fantasytalking 2.957 10.559 47.124 69.332 614.462
@@ 4.2 Experimental Setups
Evaluation Metrics. For motion-video alignment, we adopt the widely used Sync-C and Sync-D metrics chung2017out to measure the confidence and distance between visual and audio content. Hand quality and motion richness are measured using hand keypoint confidence (HKC) lin2024cyberhost and hand keypoint variance (HKV) lin2024cyberhost , respectively. These metrics validate the DPO results after integrating preference data with human motion alignment, as well as the improvements in highly dynamic motion enabled by the unified temporal motion modulation module. To evaluate facial fidelity, we employ E-FID deng2019accurate , which quantifies expression discrepancies by computing the Fréchet Inception Distance (FID) of facial expression parameters extracted via face reconstruction. For visual quality assessment, we utilize FID, FVD unterthiner2018towards , PSNR, and SSIM.
Baseline. We conduct comprehensive comparisons across two dominant architectural paradigms for audio-driven portrait animation: UNet-based and DiT-based approaches. For UNet-based methods, we benchmark against SadTalker zhang2022sadtalker , DreamTalk ma2023dreamtalk , AniPortrait wei2024aniportrait , Sonic ji2024sonic , and Hallo xu2024hallo . Our UNet variant, denoted as Ours (UNet), extends Hallo by integrating human preference data and direct preference optimization (DPO). Within the DiT-based category, we evaluate Hallo3 cui2024hallo3 and FantasyTalking wang2025fantasytalking , while our DiT implementation (Ours (DiT)) builds upon the Wan2.1 backbone wan2025 , incorporating unified temporal motion modulation and DPO-enhanced training. To validate generalization to skeletal-guided synthesis, we further compare with pose-driven half-body animation methods: AnimateAnyone hu2024animate , MimicMotion zhang2024mimicmotion , and EchoMimic-v2 meng2024echomimicv2 .
@@ 4.3 Comparison with State-of-the-Art
AnimateAnyone hu2024animate - 34.350 0.824 20.482 0.761 51.439 453.528
MimicMotion zhang2024mimicmotion - 34.783 0.846 19.186 0.713 49.039 468.004
EchoMimic-v2 meng2024echomimicv2 6.624 33.284 0.863 21.421 0.796 38.461 301.347
@@ 4.4 Ablation Study and Discussion
FantasyTalking wang2025fantasytalking 3.38 3.33 3.50 2.73
DreamTalk ma2023dreamtalk 2.83 2.88 2.68 3.20
Hallo3 cui2024hallo3 3.90 3.73 3.93 3.80
Sonic ji2024sonic 4.18 4.08 4.13 4.03